1+5+13+25+.........

Grade:Upto college level

2 Answers

Shankararama Sharma
13 years ago

The answer of sum upto n terms is,

(2n+1)(n+1)+ 3n(n+1) + 3n

Fawz Naim
13 years ago

Let

Sn= 1+5+13+25+........T(n-1)+Tn

Sn=     1+5+13+25.................T(n-1)+Tn

__________________________________________

now subtracting the two equations

0=1+4+8+12+.......upto n terms -Tn

Tn=1+(4+8+12+.......upto (n-1)terms)

Tn=1+((n-1)/2[8+(n-2)4]

solve it to get the nth term

and to solve it use summation method